A spaceship of mass 2.70 106 kg is to be accelerated to a speed of 0.695c. (a) What minimum amount of energy does this acceleration require from the spaceship's fuel, assuming perfect efficiency? J (b) How much fuel would it take to provide this much energy if all the rest energy of the fuel could be transformed to kinetic energy of the spaceship? kg
Answer #1

mo 2.70x1o*kg 0.695c 1-(0.695 m-1.390m Kīinetic Energy - Total Energy - rest mass energy m,e (1.390-1) (0.390 0.390 x 2.70x10

E-mc2 m= C2 9.477 x1022 (3x109) -10.53x10 kg m= 10.53× 105kg (cx) The mass of the fuel is therm ass increase = 0.390m 。 0.390